# Hao Sun, Weiwei Deng, Feng Sun, Furu Wei, Qi Zhang Microsoft Corporation

### v1

@@ Abstract
Document retrieval techniques form the foundation for the development of large-scale information systems. The prevailing methodology is to construct a bi-encoder and compute the semantic similarity. However, such scalar similarity is difficult to reflect enough information and impedes our comprehension of the retrieval results. In addition, this computational process mainly emphasizes the global semantics and ignores the fine-grained semantic relationship between the query and the complex text in the document. In this paper, we propose a new method called Generation Augmented Retrieval (GeAR) that incorporates well-designed fusion and decoding modules. This enables GeAR to generate the relevant text from documents based on the fused representation of the query and the document, thus learning to "focus on" the fine-grained information. Also when used as a retriever, GeAR does not add any computational burden over bi-encoders. To support the training of the new framework, we have introduced a pipeline to efficiently synthesize high-quality data by utilizing large language models. GeAR exhibits competitive retrieval and localization performance across diverse scenarios and datasets. Moreover, the qualitative analysis and the results generated by GeAR provide novel insights into the interpretation of retrieval results. The code, data, and models will be released after completing technical review to facilitate future research.
@@ 1 Introduction
Document retrieval serve as the foundational technology behind large-scale information systems, playing a crucial role in applications such as web search, open-domain question answering (QA) (Chen et al., 2017; Karpukhin et al., 2020), and retrieval-augmented generation (RAG) (Lewis et al., 2020; Liu et al., 2024a; Gao et al., 2024). The predominant approach in passage retrieval is to construct a bi-encoder model. In this architecture, queries and documents are encoded separately, transforming each into vector representations that enable computation of their semantic similarity in a high-dimensional space.
@@
However, this similarity calculation process faces several challenges. First, the complex semantic relationship between query and document is mapped to a scalar similarity, which cannot reflect enough information and is difficult to understand Brito and Iser (2023). Second, when dealing with long documents, such as those with 256, 512, or even more tokens, identifying the section most relevant to the query and contributing most to the similarity is highly desirable but challenging to achieve Luo et al. (2024); Günther et al. (2024). Moreover, many NLP tasks, such as sentence selection, search result highlighting, needle in a haystack (Liu et al., 2024b; An et al., 2024; Wang et al., 2024), and fine-grained citations (Gao et al., 2023; Zhang et al., 2024), require a deep and fine-grained understanding of the text.
Given this need for fine-grained understanding, the bi-encoder that simply aligns the entire document to the query seems insufficient, as its conventional contrastive loss mainly emphasizes global semantics Khattab and Zaharia (2020). To complement this core localization capability of the retriever, we propose a novel and challenging fundamental question: Can we enhance and integrate the information localization capability of existing retrievers without sacrificing their inherent retrieval capabilities?
@@
To address these challenges, we proposed a novel approach GeAR (Generation-Augmented Retrieval). Specifically, we construct the data into (query-document-information) triples, still using contrastive learning to optimize the similarity between the query and the document. At the same time, we design a text decoder to generate the relevant fine-grained information in the document given the query and document to enhance the retrieval and localization capabilities. Although the concept is simple, there are many challenges. First, it is difficult to find sufficient data to support our solution to this problem in previous research work. Second, the training objectives of retrieval and generation tasks, model architectures, and more design details, as well as how to effectively train the models, have not been fully explored. To this end, we explored a complete pipeline from data synthesis, structure design, to model training. Overall, our contributions are summarized as follows:
  • We proposed GeAR, which enhances the model’s ability to understand and locate text in a fine-grained manner by jointly modeling natural language understanding and natural language generation. At the same time, the inference process is very flexible to handle different tasks.
  • We abstract a new retrieval task that takes into account the problems present in the current retrieval scenario. To solve this task and to support model training, we built a pipeline to synthesize a large amount of high quality data using LLM.
  • Through extensive experiments, GeAR has shown competitive performance in retrieval tasks and fine-grained information localization tasks. At the same time, GeAR can also generate relevant information based on the query and document to help us understand the retrieval results, bringing a new perspective to the traditional retrieval process.
@@
Figure 1: Comparison of functionality between classical retriever and GeAR. GeAR is designed to handle both document retrieval and fine-grained unit localization simultaneously, while also generating auxiliary information for reference.
@@ 2.1 Embedding-based Retrieval
Embedding-based retrieval has emerged as a cornerstone of modern information retrieval systems, enabling efficient semantic search through dense vector representations. Early approaches like Word2Vec (Mikolov et al., 2013) and GloVe (Pennington et al., 2014) demonstrated the potential of learning distributed word representations, while more recent transformer-based models such as BERT (Devlin et al., 2019) have pushed the boundaries of contextual embeddings. Bi-encoder architectures (Reimers and Gurevych, 2019) have become particularly popular for retrieval tasks Huang et al. (2013). Recent advances include contrastive learning objectives (Karpukhin et al., 2020; Wang et al., 2022; Li et al., 2023; Gao et al., 2021), hard negative mining strategies (Xiong et al., 2021), and knowledge distillation techniques (Hofstätter et al., 2021) to improve embedding quality while maintaining computational efficiency. Muennighoff et al. (2024) explored how to generate text and provide excellent semantic representation by distinguishing task instructions.
Multimodal information retrieval also relies on high-quality semantic representations, where the embedding space serves to bridge different modalities, including text, images, and video. Vision language models such as CLIP (Radford et al., 2021), ALBEF (Li et al., 2021), and BLIP (Li et al., 2022) have demonstrated remarkable zero-shot capabilities by learning joint embeddings derived from large scale image-text pairs. These advances made cross-modal retrieval tasks such as text-to-image search and image-to-text retrieval possible.
2.2 Information Localization
Information localization in massive corpora and contents has become a key research direction for improving response accuracy and factual basis. The classic methods used RNN or BERT to compute token representations and trained a classifier for information extraction (Seo, 2016; Wang, 2016; Chen et al., 2017; Xu et al., 2019). The heuristic hierarchical approach involves further chunking the document and then calculating the semantic similarity with the query on the chunked sentences or units for localization. However, finer chunking also results in increased computational complexity and semantic incoherence Yang et al. (2016); Liu et al. (2021); Arivazhagan et al. (2023). With the development of generative models, there have been many recent efforts to enhance the model’s ability to find a needle in a haystack (Liu et al., 2024b; An et al., 2024; Wang et al., 2024), that is, to locate key information such as sentences in long texts. Another type of similar task is to have the model add reference information to the original text when generating responses (Gao et al., 2023; Zhang et al., 2024). Coincidentally, there have been some recent works focusing on improving the regional level understanding ability of multimodal large language models (MLLMs) (Chen et al., 2024). Despite these advances, we have found that there is currently little focus on fine-grained information localization during the retrieval stage.
@@ 3.1 Preliminaries
In this work, we formalize the retrieval task with localization as follows: Let a document corpus as $\mathbb{D}$, which contains $N$ documents $\{d_{1},...,d_{i},...,d_{N}\}$. Each of these documents $d_{i}$ contains a number of fine-grained information units $\{u_{1},...,u_{l_{i}}\}$, such as sentences, where $l_{i}$ is the units number of $d_{i}$. Our goal is to find a retrieval method $f(\cdot)$, which can retrieve the relevant document $d$ from $\mathbb{D}$, as well as the fine-grained unit $u$ from $d$ given query $q$:
@@
In this work, we explicitly define the process as two tasks, (1) the document retrieval task and (2) the fine-grained unit localization task, as Figure 1 showing. It can be seen that the triples of query, document, and unit, represented by the symbols $(q,d,u)$, are fundamental to the definition and resolution of this task.
@@ 3.2 Data construction
In this work, we focus on two common retrieval scenarios: (1) Question Answer Retrieval (QAR) and (2) Relevant Information Retrieval (RIR). In the following sections, we will introduce how the data are constructed and how they correspond to the triples $(q,d,u)$ mentioned above.
Question Answer Retrieval  In this scenario, the query $q$ is in the form of a question, and the goal is to retrieve reference documents $d$ that support answering the question and fine-grained sentences $u$ that contain the answer.
@@
Relevant Information Retrieval  In this scenario, the query $q$ is in the form of a few phrases or keywords, the objective is to retrieve the documents $d$ that correspond to the query and the fine-grained sentences $u$ in the documents that are most relevant to the query. The scenario is very close to what users normally do when they search for information on the web. The challenge is that we have difficulty in finding suitable data in the current public dataset to drive our problem solving. Therefore, we constructed a pipeline to synthesize high quality data using a large language model. Specifically, we selected high quality Wikipedia documents (Foundation, ), from which we will sample sentences of appropriate length and whose subject is not a pronoun as $u$. Then we will leverage LLM to rewrite them as queries $q$. After de-duplication and relevance filtering, we get promising 5.8M triples. Kindly refer to Appendix A for details on complete data processing procedure.
@@ 3.3 Model Structure
This section presents the architecture of GeAR. It is our intention that the model not only has powerful retrieval capability, but also has the ability to locate key information in documents. Inspired by advances in multimodal representation learning (Li et al., 2021, 2022; He et al., 2020), we revisit the task from a modal alignment perspective. Documents and queries can be considered as two modalities. We facilitate semantic alignment between documents and queries via a bi-encoder, and enable the model to learn to attend to fine-grained query-related information in the document via a fusion encoder and a generation task. The overview of the GeAR structure is illustrated in Figure 2.
@@
Fusion Encoder The fusion encoder share most of the parameters with query encoder, but have an extra learnable cross attention module. In this part, the document embeddings from $E_{d}(\cdot)$ are fused with the query embeddings through cross attention at each layer of the fusion encoder.
Text Decoder The text decoder receives the fusion embeddings and generates fine-grained information111 Note that in the generation task of the QAR scenario, the ground truth is the answer itself, not the sentence $u$. But in the RIR scenario and the localization task, we all used the sentence $u$. in the document based on the given query and document. It uses a unidirectional causal attention instead of a bidirectional self-attention. A specific [Decode] token is added to identify the beginning of the sequence. The subsequent auto-regressive decoding process will interact with the generated tokens and fusion embeddings to generate text.
@@ 3.4 Training Objectives
In this section, we present the training objectives of GeAR. We make the model capable of both retrieval as well as fine-grained semantic understanding and localization through joint natural language understanding and natural language generation modeling.
@@
Language Modeling Loss (LM)  The introduction of LM loss is key to enhancing the information localization capability of GeAR. LM activates the text decoder, which enables the model to generate relevant information using the fusion embeddings of document and query. It guides the model to learn the fine-grained semantic fusion of query and document. LM optimizes the cross entropy loss over the entire vocabulary, maximizing the likelihood of the ground truth text. The overall loss of GeAR is the sum of $\mathcal{L}_{\mathrm{CL}}$ and $\mathcal{L}_{\mathrm{LM}}$:
$\mathcal{L}_{\mathrm{GeAR}}=\mathcal{L}_{\mathrm{CL}}+\mathcal{L}_{\mathrm{LM}}$ (3)
@@ 3.5 Inference
GeAR’s inference process is diverse and flexible. In this section, we introduce various usages of GeAR to accomplish different tasks.
Documents Retrieval For this task, we can use the bi-encoder part of GeAR to compute the similarity between query and document like the previous classic retrieval method, without introducing any additional parameters and computation cost.
Fine-Grained Units Localization The fusion encoder in GeAR calculates the fusion embedding of query and document through cross attention. We use the cross attention weights of the query on the tokens in the document to locate the units that the query pays the most attention to in the document.
Information Generation For this task, we feed the fusion embedding to the text decoder and enable autoregressive decoding. In GeAR, information generation is actually an auxiliary task, and we will present the generative performance of the model in experiments, both in terms of quantitative metrics and qualitative analysis.
@@ 4 Experiments
In this section, we first introduce the experimental setup, and then we show the overall performance of each task and more detailed analysis experiments.
@@ 4.1 Setup
Datasets For Question Answer Retrieval, we sampled 30M data from PAQ (Lewis et al., 2021) datasets to train GeAR, and sampled 1M documents and 20k queries as test set. We also evaluate the performance on another 3 QA datasets: SQuAD (Rajpurkar et al., 2016), NQ (Kwiatkowski et al., 2019), and TriviaQA (Joshi et al., 2017). These test datasets are all held out to observe the generalizability of compared methods. For Relevant Information Retrieval, we leverage the synthesized 5.8M data, of which 95% is used for training and 5% is reserved for the test set. Specific dataset statistics are in Appendix B.
Training Details To better observe the effectiveness of GeAR, we use "BERT-base-uncased" (Devlin et al., 2019) to initialize the encoders in GeAR. We trained the model for 10 epochs using a batch size of 48 (QAR) / 16 (RIR) on 16 AMD MI200 GPUs with 64GB memory. We use the AdamW (Loshchilov, 2017) optimizer with a weight decay of 0.05. The full hyperparameters and training settings are detailed in Appendix C.
Baselines We compare our approach to two classes of baseline methods, one class of text representation models that have been adequately trained on a large corpus, including SBERT (Reimers and Gurevych, 2019), specifically "all-mpnet-base" (Song et al., 2020), E5 (Wang et al., 2022), BGE (Xiao et al., 2024), and GTE (Li et al., 2023). We use both base-level models for this comparison. The other category consists different training pipelines that unify the training data and starting points, including SBERT (Reimers and Gurevych, 2019) and BGE (Xiao et al., 2024). We retrained them all using the "bert-base-uncased" to initialize and aligned the training data, referred to as SBERTRT and BGERT in the following.
SQuAD NQ TriviaQA PAQ RIR

### v2

@@ Abstract
Document retrieval techniques are essential for developing large-scale information systems. The common approach involves using a bi-encoder to compute the semantic similarity between a query and documents. However, the scalar similarity often fail to reflect enough information, hindering the interpretation of retrieval results. In addition, this process primarily focuses on global semantics, overlooking the finer-grained semantic relationships between the query and the document’s content. In this paper, we introduce a novel method, Generation Augmented Retrieval (GeAR), which not only improves the global document-query similarity through contrastive learning, but also integrates well-designed fusion and decoding modules. This enables GeAR to generate relevant context within the documents based on a given query, facilitating learning to retrieve local fine-grained information. Furthermore, when used as a retriever, GeAR does not incur any additional computational cost over bi-encoders. GeAR exhibits competitive retrieval performance across diverse scenarios and tasks. Moreover, qualitative analysis and the results generated by GeAR provide novel insights into the interpretation of retrieval results. The code, data, and models will be released at https://github.com/microsoft/LMOps.
@@ 1 Introduction
Document retrieval serve as the foundational technology behind large-scale information systems, playing a crucial role in applications such as web search, open-domain question answering (QA) (Chen et al., 2017; Karpukhin et al., 2020), and retrieval-augmented generation (RAG) (Lewis et al., 2020; Liu et al., 2024a; Gao et al., 2024). The predominant approach in passage retrieval is to construct a bi-encoder model (Reimers and Gurevych, 2019). In this framework, queries and documents are encoded separately, converting each into vector representations that enable computation of their semantic similarity in a high-dimensional space.
@@
However, this similarity calculation process faces several challenges. First, the complex semantic relationship between query and document is mapped to a scalar similarity, which cannot reflect enough information and is difficult to understand Brito and Iser (2023). Second, when dealing with long documents, such as those with 256, 512, or even more tokens, identifying the section most relevant to the query and contributing most to the similarity is highly desirable but challenging to achieve Luo et al. (2024); Günther et al. (2024). Moreover, many NLP tasks, such as sentence selection, search result highlighting, needle in a haystack (Liu et al., 2024b; An et al., 2024; Wang et al., 2024), and fine-grained citations (Gao et al., 2023; Zhang et al., 2024), require a deep and fine-grained understanding of the text. Given this need for fine-grained understanding, the bi-encoder that simply aligns the full document to the query seems insufficient, as its conventional contrastive loss mainly emphasizes global semantics Khattab and Zaharia (2020). To complement this core capability of the retriever, we propose a novel and challenging fundamental question: How to make the retriever have both global and local understanding and retrieval capabilities?
@@
Although the concept is intutive, several challenges remain. First, it is difficult to construct sufficient data to support effective solutions to this problem in previous research work. Second, the training objectives, model architectures, design details, as well as how to effectively train the models, have not been fully explored. To address these challenges, we propose a novel approach GeAR (Generation-Augmented Retrieval). In it, we build a pipeline to efficiently synthesize large amounts of high-quality (query-document-information) triples by utilizing large language models. In terms of method, GeAR retains to leverage contrastive learning to optimize the similarity between the query and the global document. To improve the interaction between local information and queries, we design a text decoder that generates fine-grained information from the document in response to a given query. This enhances the model’s ability to understand local semantics. In this way, GeAR can handle both the retrieval of global documents and local information simultaneously.
We conduct extensive experiments on two retrieval tasks, and compared with the BGE and BGE-Reranker-L, GeAR achieves 3.5% and 12.9% relative improvements on global document retrieval and local information retrieval tasks respectively. GeAR’s versatility and visual analysis also shed new light on the interpretability and comprehensibility of retrieval results.
Overall, our contributions are summarized as follows:
  • We introduce a new global-local retrieval task, which presents challenges for both document retrieval and fine-grained information retrieval within documents.
  • We introduce GeAR, which augmented the model’s global and local understanding and retrieval capabilities of documents by incorporating a generation task.
  • Through extensive experiments, GeAR has shown competitive performance across various retrieval tasks. GeAR’s versatility also makes the retrieval results more explainable.
@@
Figure 1: Comparison of functionality between classical retriever and GeAR. GeAR is designed to handle both global document retrieval and local information retrieval simultaneously. In addition, GeAR can generate information based on the query for reference.
@@ 2.1 Embedding-based Retrieval
Embedding-based retrieval has emerged as a cornerstone of modern information retrieval systems, enabling efficient semantic search through dense vector representations. Early approaches like Word2Vec (Mikolov et al., 2013) and GloVe (Pennington et al., 2014) demonstrated the potential of learning distributed word representations, while more recent transformer-based models such as BERT (Devlin et al., 2019) have pushed the boundaries of contextual embeddings. Bi-encoder architectures (Reimers and Gurevych, 2019) have become particularly popular for retrieval tasks Huang et al. (2013). Recent advances include contrastive learning objectives (Karpukhin et al., 2020; Wang et al., 2022; Li et al., 2023; Gao et al., 2021) and hard negative mining strategies (Xiong et al., 2021) to improve embedding quality. Muennighoff et al. (2024) explored how to generate text and provide excellent semantic representation by distinguishing task instructions. Multimodal information retrieval also relies on high-quality semantic representations, where the embedding space serves to bridge different modalities, including text, images, and video. Vision language models such as CLIP (Radford et al., 2021), ALBEF (Li et al., 2021), and BLIP (Li et al., 2022) have demonstrated remarkable zero-shot capabilities by learning joint embeddings derived from large scale image-text pairs.
2.2 Fine-grained Information Mining
Mining fine-grained information in a long context during retrieval has become a key challenge for efficient information retrieval. The naive heuristic hierarchical approach involves further chunking documents and then calculate semantic similarity with the query on the chunked sentences. However, finer chunking easily leads to increased computational complexity and semantic incoherence Yang et al. (2016); Liu et al. (2021); Arivazhagan et al. (2023). In question-answering tasks, RNN or BERT is often used to compute token representations and train classifiers for information extraction (Seo, 2016; Wang, 2016; Chen et al., 2017; Xu et al., 2019). With the development of generative models, there have been many efforts to enhance the model’s ability to find a needle in a haystack (Liu et al., 2024b; An et al., 2024; Wang et al., 2024). Another similar task is to have the model add reference information to the original text when generating responses (Gao et al., 2023; Zhang et al., 2024). Coincidentally, some recent research is dedicated to improving the region-level understanding ability of multimodal large language models (MLLMs) (Chen et al., 2024).
Despite these advances, we find that these works often rely on heavy decoder-only models that are independent of the retrieval model, but few focus on mining fine-grained information during the retrieval stage.
@@ 3.1 Preliminaries
In this work, we formalize the global-local retrieval task as follows: Let a document corpus as $\mathbb{D}$, which contains $N$ documents $\{d_{1},...,d_{i},...,d_{N}\}$. Each of these documents $d_{i}$ contains a number of fine-grained information units $\{u_{1},...,u_{l_{i}}\}$, such as sentences, where $l_{i}$ is the units number of $d_{i}$. Our goal is to find a retrieval method $f(\cdot)$, which can retrieve the relevant document $d$ from $\mathbb{D}$, as well as the fine-grained information $u$ from $d$ given query $q$:
@@
In this work, we explicitly define the process as two tasks, (1) the global document retrieval and (2) the local information retrieval, as shown in Figure 1.
@@ 3.2 Data construction
In this work, we consider two main retrieval scenarios: Question Answer Retrieval (QAR) and Relevant Information Retrieval (RIR). In the following sections, we introduce how the data is constructed and outline the specific goals of the retrieval tasks in each scenario.
Question Answer Retrieval  In this scenario, the query $q$ is in the form of a question, and the goal is to retrieve (1) the reference documents $d$ that support answering the question and (2) the fine-grained sentences $u$ that contain the answer.
@@
Relevant Information Retrieval  This scenario closely mirrors typical user behavior when searching for information on search engines. The query $q$ is typically a few phrases or keywords, the objective is to retrieve (1) the documents $d$ that correspond to the query and (2) the fine-grained sentences $u$ in the documents that are most relevant to the query. However, a significant challenge in this scenario is the difficulty of collecting suitable data from existing public datasets to address this problem. To overcome this, we construct a pipeline to synthesize high quality data using a large language model. Specifically, we select high quality Wikipedia documents (Foundation, ), from which we sampled sentences of appropriate length and whose subject is not a pronoun as $u$. Then we leverage LLM to rewrite these sentences as queries $q$. After applying de-duplication and relevance filtering, we obtain a promising set of 5.8M triples. Kindly refer to Appendix A for details on complete data processing procedure.
@@ 3.3 Model Structure
This section introduces the architecture of GeAR. It is our intention to enable the model to have both global and local text retrieval capabilities. Inspired by advances in multimodal representation learning (Li et al., 2021, 2022; He et al., 2020), we revisit the task from the perspective of modality alignment. Documents and queries can be regarded as two modalities. We facilitate semantic alignment between documents and queries via a bi-encoder, and enable the model to learn to focus on fine-grained query-related information in documents via a fusion encoder and a generation task. The overview of the GeAR structure is illustrated in Figure 2.
@@
Fusion Encoder The fusion encoder share most of the parameters with query encoder, but have an lightweight learnable cross attention module. In this part, the document embeddings from $E_{d}(\cdot)$ are fused with the query embeddings through cross attention at each layer of the fusion encoder.
Text Decoder The text decoder receives the fusion embeddings and generates fine-grained information111 Note that in the QAR scenario, the ground truth for the generation is the answer itself, not the full sentence $u$ in which answer appears. in the document based on the given query and document. It uses a unidirectional causal attention instead of a bidirectional self-attention. A specific [Decode] token is added to identify the beginning of the sequence. The subsequent auto-regressive decoding process will interact with the generated tokens and fusion embeddings to generate text.
@@ 3.4 Training Objectives
In this section, we introduce the training objectives of GeAR. Through the joint modeling of natural language understanding and natural language generation, GeAR can handle global document retrieval and local information retrieval simultaneously.
@@
Language Modeling Loss (LM)  The introduction of LM loss is crucial for enhancing the local information retrieval capability of GeAR. LM activates the text decoder, enabling the model to generate relevant intrinsic information by leveraging the fusion embeddings of document and query. It guides the model to learn the fine-grained semantic fusion between query and document. LM optimizes the cross-entropy loss over the entire vocabulary, maximizing the likelihood of the ground truth text. The overall loss of GeAR is the sum of $\mathcal{L}_{\mathrm{CL}}$ and $\mathcal{L}_{\mathrm{LM}}$ with a optional weight $\alpha$:
$\mathcal{L}_{\mathrm{GeAR}}=\mathcal{L}_{\mathrm{CL}}+\alpha*\mathcal{L}_{\mathrm{LM}}$ (3)
@@ 3.5 Inference
GeAR’s inference process is flexible. In this section, we introduce various usages of GeAR to accomplish different tasks.
Global Documents Retrieval For this task, we can use the bi-encoder part of GeAR to compute the similarity between query and document like the previous classic retrieval method, without introducing any additional parameters and computation cost.
Local Information Retrieval The fusion encoder in GeAR interacts query and document via cross attention. The cross attention weights between each sentence in the document and the query reflect which information the model prioritizes. We rank the sentences based on these weights to retrieve the most relevant fine-grained information from the document.
@@ 4 Experiments
In this section, we first outline the experimental setup, and then we discuss the overall performance of each task and a more detailed analysis.
@@ 4.1 Setup
Datasets For Question Answer Retrieval, we sampled 30M data from PAQ (Lewis et al., 2021) datasets to train GeAR, and sampled 1M documents and 20k queries as the test set. To verify the generalization ability of methods, we also evaluate the performance on three additional held-out datasets: SQuAD (Rajpurkar et al., 2016), NQ (Kwiatkowski et al., 2019), and TriviaQA (Joshi et al., 2017). For Relevant Information Retrieval, we leverage the synthesized 5.8M data, of which 95% is used for training and 5% is reserved for the test set. Specific dataset statistics are in Appendix B.
Training Details "bert-base-uncased" (Devlin et al., 2019) is used to initialize the encoders in GeAR. The decoder also has 110M parameters, but is randomly initialized. We train GeAR for 10 epochs using batch size of 48 (QAR) / 16 (RIR) on 16 AMD MI200 GPUs. We set the weight $\alpha=0.25$. We use the AdamW (Loshchilov, 2017) optimizer with a weight decay of 0.05. The full hyperparameters and training settings are detailed in Appendix C.
Baselines We compare GeAR with two types of baselines, one is the text embedding models that have been adequately pre-trained on a large corpus, including SBERT (Reimers and Gurevych, 2019), E5 (Wang et al., 2022), BGE (Xiao et al., 2024), GTE (Li et al., 2023) and ColBERT-QA (Khattab et al., 2021). The models involved in the comparison are all base versions. Since the training data of the pre-trained model partially overlaps with the evaluation data, their performance are used as an important reference. To ensure a fairer comparison, we retrain SBERT [^2] https://huggingface.co/sentence-transformers/all-mpnet-base-v2 (Reimers and Gurevych, 2019) and BGE [^3] https://github.com/FlagOpen/FlagEmbedding (Xiao et al., 2024) using the open sourced training pipelines with aligned training data and initialization, referred to as SBERTRT and BGERT in the following. In addition, we also compare with the more complex BGE reranker Base and Large (Xiao et al., 2024) in the Local Information Retrieval task. In the section 4.2, we underline the best performance of the pre-trained models and bold the best performance of the retrained models.
4.2 Overall performance
In this section, we present the overall performance on global document retrieval and local information retrieval.
Global Documents Retrieval Firstly, Table 2 reports the comparison with existing methods on global documents retrieval task. We find that GeAR delivers competitive performance across multiple datasets even with only tens of millions of training data, demonstrating efficient data utilization. As a reference, the pre-trained SBERT model used 1.17B sentence pairs. GeAR achieves the state-of-the-art performance on the three datasets SQuAD, PAQ, and RIR, and is slightly weaker than the pre-trained GTE on the NQ dataset. It only lags significantly behind on TriviaQA, but is also better than ColBERT-QA and E5. Compared with ColBERT, GeAR introduces a generation task to explicitly model the alignment relationship between queries and fine-grained semantic fragments of documents, which not only improves the retrieval performance but also reduces the delayed interaction and the increase in space complexity caused by storing multiple vectors. At the same time, GeAR outperforms the retrained model in all metrics. Compared with BGERT, GeAR achieves a relative improvement of 3.5% in average Recall@5, highlighting the effectiveness of our training method. In Section 4.3, we further discuss the role of the generation task and its effect on model performance.
Method SQuAD NQ TriviaQA PAQ RIR